OpenStudy (anonymous):

Graph the following inequality and then select the correct graph below. x + 2 ≥ 1

OpenStudy (anonymous):

 am i correct?

HanAkoSolo (jamierox4ev3r):

which line is which? i mean which is the axis?

OpenStudy (anonymous):

Solve for x.

OpenStudy (anonymous):

the one to the left is the axis

OpenStudy (anonymous):

x=y+1

OpenStudy (anonymous):

The graph is incorrect. Solve you equation for x and graph again.

HanAkoSolo (jamierox4ev3r):

x=-1, so the line to the left should be the line for the equation THe reason why is because: x+2\(\ge\)1 -2 -2 ----------- x\(\ge\)-1

OpenStudy (anonymous):

Which equation are you graphing?

OpenStudy (anonymous):

 how about now?

HanAkoSolo (jamierox4ev3r):

now that is correct good job :)

OpenStudy (anonymous):

ok thank you

HanAkoSolo (jamierox4ev3r):

but make sure the line is correct lol :P

OpenStudy (anonymous):

Less than or equal to is not just a line, it is always a shaded area.